OpenStudy (anonymous):

show that 11^n -4^n is divisible by 7

OpenStudy (isaiah.feynman):

What could n mean?

OpenStudy (zzr0ck3r):

this is induction, n is natural

OpenStudy (isaiah.feynman):

I don't think that would work anyway, 7 is a prime. It won't divide any number perfectly.

OpenStudy (anonymous):

zzr0ck3r is right except I haven't got a clue what to do

OpenStudy (anonymous):

seems to work for n=1 and 2 so I assume it works for the rest of natural numbers

OpenStudy (zzr0ck3r):

so for sure this is true when n = 1 so suppose 7 divides \(11^n-4^n\) for some \(n\in \mathbb{N}\) so \(11^{n+1}-4^{n+1} = 11*11^n-4*4^n = (11^n-4^n)+7*11^n+3*11^n-3*4^n=\\(11^n-4^n)+7*11^n+3(11^n-4^n) = 4(11^n-4^n)+7*11^n=4*7k+7*11^n\) surely 7 divides this number so by induction we are done.

OpenStudy (isaiah.feynman):

I really need to teach myself this topic throughly! Know any good books?

OpenStudy (zzr0ck3r):

I would just google "induction". Its really not that bad. just think of dominoes. Any standard group theory or number theory book will be sufficient.

OpenStudy (anonymous):

I slept on it and that is what was in my head this morning lol....the answer you gave